7.0 g of propane C3H8 and an abundance of O2 how many grams of water can be produced
1
Expert's answer
2020-04-29T13:28:59-0400

Solution.

C3H8+5O2=3CO2+4H2O.

n(C3H8)=m/M=7/44=0,159 mole;

Calculate the chemical amount of water from the reaction (look at the coefficients):

n(H2O)=(0,159*4)/1=0,636 mole;

m(H2O)=n*M=0,636*18=11,448 g.

Answer: 11,448 g of water.